OpenStudy (sf_r):

Find the equation, in vertex form, for a parabola with a vertex of (2,-5) that goes through the point (3,4).

OpenStudy (sf_r):

@ganeshie8

OpenStudy (sf_r):

@ILovePuppiesLol

OpenStudy (sf_r):

@AloneS

OpenStudy (sf_r):

@mathmath333

OpenStudy (sf_r):

@jameshorton

OpenStudy (sf_r):

@dayakar

OpenStudy (sf_r):

I have the answer choices to the question if you need them

OpenStudy (sf_r):

@ParthKohli

OpenStudy (dayakar):

Standard form of equation for a parabola: y=A(x-h)^2+k, with (h,k) being the (x,y) coordinates of the vertex

OpenStudy (sf_r):

So y=a(x-2)^2-5

OpenStudy (dayakar):

(x,y) = (3,4) , (h,k) =(2,-5) 4 = A (3-2) ^2 + (-5) simplify and find A

OpenStudy (sf_r):

okay

OpenStudy (sf_r):

So first i do the square, right?